300 j of heat added to a system and 250 j of work is added by the system. what is the change in internal energy by the system?


1
Expert's answer
2022-03-28T23:54:03-0400

From the first law of thermodynamics we know that,

ΔU=q+W

In this case, q = 300 J, W = 250 J

∴ΔU=300+ 250= 550J